OpenStudy (anonymous):

please help I'm willing to give medals and I want to be a fan What is $\frac{1}{\frac{1}{7}\cdot\frac{1}{-8}}$?

OpenStudy (agreene):

\[\frac1{\frac17\cdot\frac{-1}8}\] is this the question?

OpenStudy (anonymous):

ummm almost, the 8 is supposed to be -8 and the 1 is just 1

OpenStudy (agreene):

well that doesnt matter, really. you can put that - wherever as long as it's attached to the fraction. so: \[\frac{1}{\frac17\cdot\frac1{-8}}=-\frac{1}{\frac17\cdot\frac1{8}}=-\frac1{\frac1{56}}=-56\]

OpenStudy (anonymous):

oh I see